how is this solved I am unsure how to proceed
4/9-1/9

2/y-3+3/3-x

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

There are no equal signs so they can't be "solved".
Second, one problem at a time.
4/9-1/9= (4-1)/9=3/9=1/3

Answer by stanbon(75887)

how is this solved I am unsure how to proceed
4/9-1/9

2/y-3+3/3-x
================
There is nothing to solve; you have no equations.
------------------------------------------
Cheers,
Stan H.